OpenStudy (anonymous):

Multivariate Calc question: Find all relative extrema of x^2y^2 subject to the constraint 4x^2 + y^2 = 8 Thanks in advance!

OpenStudy (anonymous):

I think: Critical points in the interior are when both partials=0 (or one doesn't exist). This means that the only interior critical point is (0,0)

OpenStudy (anonymous):

Or is this a lagrange multipliers question?

OpenStudy (anonymous):

If it is Lagrange, would the system be: \[<2y ^{2} x, 2x ^{2}y> = \lambda<8x, 2y>\] ?

OpenStudy (anonymous):

Yeah, I'm pretty lost on this one.

OpenStudy (turingtest):

hm... (0,0) is not within the constraint 4x^2 + y^2 = 8, so it's out of bounds I do think it's that kind of problem though, not lagrange

OpenStudy (turingtest):

it is an equals sign in the constraint, not less-than or equal-to, right?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

That's what made me think Lagrange after I realized that (0,0) isn't in the domain.

OpenStudy (anonymous):

Ooh. We know x and y can't both be 0 because that isn't in the domain, so you can cancel terms in the Lagrange equation, yielding: 2y^2 = 8 and 2x^2 = 2 along with 4x^2 + y^2 = 8

OpenStudy (anonymous):

y = 2 and x = 1 according to that, so z=4

OpenStudy (anonymous):

or -1 and -2, which also yields 4

OpenStudy (turingtest):

yes, I came to the same conclusion by doing a simple substitution lagrange is unnecessary I guess

OpenStudy (anonymous):

Although that doesn't yield any minima... which I can't imagine is true given those equations

OpenStudy (turingtest):

I think you were right actually, it is a Lagrange problem... in which case I can't really help since I'm not so good at it I'll have to read up on it, but hopefully I can answer it next time

OpenStudy (anonymous):

Great, well thanks anyway. I really appreciate the help!